OpenStudy (nani200):

find all solutions to the equation cos x = -1/2

OpenStudy (welshfella):

Hint: the cosine is negative in the second and third quadrants. and the angle whose cosine is 1/2 is pi/3 radians.

OpenStudy (nani200):

i don't know how to solve it

OpenStudy (nani200):

pi/3 is in the first quadrant though

OpenStudy (welshfella):

yes - but from this you should find the 2 angles in the second and third quadrants.

OpenStudy (welshfella):



OpenStudy (welshfella):

angle in 2nd quadrant = pi - pi/3 = 2pi/3

OpenStudy (nani200):

so are the other solutions 4pi/3 and 5pi/3?

OpenStudy (welshfella):



OpenStudy (welshfella):

there is only one more solution - the one in 3rd quadrant

OpenStudy (welshfella):

pi + pi/3

OpenStudy (nani200):

so its 4pi/3

OpenStudy (welshfella):

Yes I am assuming the question only wants solutions in the range 0 to 2pi

OpenStudy (nani200):

yes

OpenStudy (welshfella):

OK so these are your 2 solutions 2pi/3 , 4pi/3 or 120 , 240 degrees

OpenStudy (nani200):

is this answer choice correct \[\left\{ \frac{ 2\pi }{ 3 } + 2n \pi, \frac{ 4\pi }{ 3 }+ 2n \pi n= 0, +/-1, +/-2,...\right\}\]